OpenStudy (anonymous):

Solve the quadratic equation: 2x2 + 11x − 6 = 0

OpenStudy (anonymous):

ANSWERS:

OpenStudy (anonymous):

use the super quadratic formula. let a=2,b=11, and c=-6

OpenStudy (anonymous):

2x^2 + 11x - 6 = 0 a = 2 b = 11 c = -6 quadratic formula: x = -b +/- sqrt b^2 - 4ac / 2a x = -11 +/- sqrt 11^2 -4(2)(-6) / 2(2) x = -11 +/- sqrt 121 + 48 / 4 x = -11 +/- sqrt 169 / 4 x = -11 +/- 13 / 4 So there you have your two equations which you'll need to do for your answer. -11 + 13 / 4 2 / 4 1/2 is your first one -11 - 13 / 4 -24/4 -6 is your second one

OpenStudy (anonymous):

So your answer is the third.

OpenStudy (anonymous):

Thanks guys !